Cold Realty Corp

Cold Realty Corp., a closely-held corporation, has an authorized capital stock of 100,000 shares with par value of P1.00/share. Of the 100,000,000 authorized shares, 25,000,000 thereof is subscribed and fully paid up by the following stockholders:

Mr. Malamig                                5,000,000

Mrs. Maginaw                            5,000,000

Mr. Tugnaw                                5,000,000

Mr. Cool                                    5,000,000

Mr. Fresh                                5,000,000

Total Shares Outstanding    25,000,000


On March 2018, Cold Realty Corp., finally decides to conduct an IPO and initially offers

25,000,000 of its unissued shares to the investing public for P1.50/share.

At the IPO, one of its existing stockholders, Mrs. Maginaw, has likewise decided to sell her entire 5,000,000 shares to the public for P1.50/share.


1. How much is the percentage tax due on the primary offering?

2. How much is the percentage tax due on the secondary offering?

3. If in June 2018, Cold Realty Corp. again decides to increase capitalization by offering

another 30,000,000 of unissued shares to the public at P2.00/share, how much is the

percentage tax due?

4. If in July 2018, Mr. Malamig decides to sell his entire stock ownership to the public at

P2.00/share, how much is the percentage tax due? 
Answer #1

1. 375K

25M+25M = 50M

25M/50M = 50%>33 1/3%


25M*1.5 = 37.5M*1% = P375K


2. 300K

5M/50M = 10%


5M*1.5 = 7.5M*4% = P300K



3. 600K

30M+50M = 80M

30M/80M = 37.5%


30M*2 = 60M*1% = P600K



4. 400K

5M/80M = 6.25%


5M*2 = 10M*4% = P400K


answered by: anonymous